OpenStudy (anonymous):

the relation a^n|b^n implies that a|b. I know its true but i don't know how to go about writing it out in proof form

OpenStudy (aum):

I suppose proof by induction.

ganeshie8 (ganeshie8):

\(a^n | b^n \implies b^n = a^nk\) for some \(k\) \(\large \implies \frac{b^n}{a^n} - k = 0 \) \(\large \implies \left(\frac{b}{a}\right)^n - k = 0 \) Clearly b/a is a zero of above polynomial, that means \(\large a|b\) since the only possible rational zeroes of polynomial \(\large x^n-k = 0\) are integers

ganeshie8 (ganeshie8):

induction gives a nice proof too, try it :)

OpenStudy (anonymous):

thank you

ganeshie8 (ganeshie8):

For induction : base case : n=1 is true since a^1 | b^1 => a|b assuming a^k|b^k => a|b, you need to prove a^k+1|b^k+1 => a|b: \(\large a^{k+1}|b^{k+1} \implies aa^k |bb^k \implies a | bk \implies a|b \)

ganeshie8 (ganeshie8):

you need to justify each step a bit thoroughly if you're doing this proof as homework

ganeshie8 (ganeshie8):

fixed a typo in last line : \(\large a^{k+1}|b^{k+1} \implies aa^k |bb^k \implies a | b\color{Red}{t} \implies a|b\)

OpenStudy (anonymous):

thank you for the help! it makes much more sense now